PLEASE HELP!!! I need help with the following question. I am using the formula PMT= iFV/(1+i)^n. However, I can't get to the final answer which I feel is $2096.
Mary and Joe would like to save up $10,000 by the end of three years from now to buy new furniture for their home. They currently have $2500 in a savings account set aside for the furniture. They would like to make equal year end deposits to this savings account to pay for the furniture when they purchase it three years from now. Assuming that this account pays 8% interest, how much should the year end payments be?
Show all work for each assignment and explain each step carefully.
Thanks in advance - lj

Answer by stanbon(75887)

I am using the formula PMT= iFV/(1+i)^n. However, I can't get to the final answer which I feel is $2096.
Mary and Joe would like to save up $10,000 by the end of three years from now to buy new furniture for their home. They currently have $2500 in a savings account set aside for the furniture. They would like to make equal year end deposits to this savings account to pay for the furniture when they purchase it three years from now. Assuming that this account pays 8% interest, how much should the year end payments be?
--------------------------------
PMT= iFV/(1+i)^n
I think the formula should be PMT = iFV/[(1+i)^n -1]
FV = $10,000-2500 = $7500
n = 3
i = 0.08
-------------
The $2500 that is in the account in the beginning will grow at a compounded
rate of 8% yearly to have a value of 2500(1+0.08)^3 = 3149.28
So the account has to gain an increased value of 10000-3149.28= $6850.72
Therefore FV = $6850.72
PMT = 0.08*6850.72/[(1+0.08)^3 - 1]
PMT = 548.06 / [1.259712 -1]
PMT = $2110.25
=====================
Cheers,
Stan H.